40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Oct 29 15:34:36 2010, Zach & Koji, Laser, GYRO, CCW loop characterization bart-simpson-generator.png
    Reply  Sat Oct 30 20:02:01 2010, Koji, Laser, GYRO, gyro characterization Oct 27, 2010 (1) open_loop_primary.pdf
       Reply  Sat Oct 30 20:18:06 2010, Koji, Laser, GYRO, gyro characterization Oct 27, 2010 (2) error_primary.pdffeedback_primary.pdf
          Reply  Sat Oct 30 20:57:19 2010, Koji, Laser, GYRO, gyro characterization Oct 27, 2010 (3) error_secondary.pdf
             Reply  Sat Oct 30 22:49:04 2010, Koji, Laser, GYRO, gyro characterization Oct 27, 2010 (4) 
                Reply  Sat Oct 30 23:41:31 2010, Koji, Laser, GYRO, gyro characterization Oct 27, 2010 (5) 
Message ID: 1120     Entry time: Sat Oct 30 22:49:04 2010     In reply to: 1119     Reply to this: 1121
Author: Koji 
Type: Laser 
Category: GYRO 
Subject: gyro characterization Oct 27, 2010 (4) 

[Zach / Koji]

Calibration of the laser PZT drive

Method:

- Lock the primary cavity.
- Inject the sinusoidal signal from DS345 to the PZT sweep input. The frequency was 1kHz. The amplitude display showed 10mVpp(@50Ohm).
- Look at the peak in the PZT feedback and the error signal.
- Also look at the error of the secondary loop. (*)

- Disconnect VCO feedback and connect DS345 to it. This gives us the calibration signal with known amplitude.
- Look at the error signal of the secondary loop. Compare it with the above (*)

Result:

The result of the signal injection into the primary loop

- The primary loop error showed the peak amplitude of 685.2410 uVrms
- The primary loop feedback showed the peak amplitude of 98.7438 uVrms
- The secondary loop error showed the peak amplitude of 64.9701 uVrms

We put the same signal into the VCO control. (Note: It is not necessary to be the same amplitude.)

- The secondary loop error showed the peak amplitude of 247.258 uVrms

Strangely the primary error signal also showed 0.808uVrms peak, which should not appear in the primary error.

Thought:

Quick calculation

- The response of the VCO driver is 100kHz/V. Assume this has 50Ohm input (TO BE CONFIRMED). The VCO drive has 10k input impedance. => 10K/(10K+50) = 0.995
- This means we shook the VCO frequency by 100kHz/V * 10mVpp * 2 * 0.995 = 2.0kHzpp
- The actual AOM changes the laser frequency twice because of its double pass configuration. ==> 4.0kHzpp ...(a)

- (a) yield 247uVrms, while the PZT injection did 685uVrms 64.97uVrms. ==> 2.8 times larger. 4.0kHz x 0.26 = 1.04kHzpp ...(b)

- PZT actuation of 99uVrms (=280uVpp) caused this (b). ==> 3.7MHz/V

To Do:

  • Confirm the above logic
    • How much was the VCO frequency actually swung?
      Calibrate the VCO driver again by putting the same signal into the VCO and look at the VCO output with the RF analyzer.
  • Calibrate the primary and secondary error signals. How much V/Hz do they have?
  • What cause the strange signal coupling from the secondary cavity to the primary cavity?
    • Is this optical, or electorical?
    • Shake the VCO freq and observe the error signal of the primary. Block the beam of the secondary at various places.
      (i.e. just before the faraday, just after the AOM, before the AOM etc...) Try to understand what is happening?
    • Is this coupling harmful?

 

ELOG V3.1.3-